# Can Direct Latent Model Learning Solve Linear Quadratic Gaussian Control?

### v2

@@ Abstract
We study the task of learning state representations from potentially high-dimensional observations, with the goal of controlling an unknown partially observable system. We pursue a direct latent model learning approach, where a dynamic model in some latent state space is learned by predicting quantities directly related to planning (e.g., costs) without reconstructing the observations. In particular, we focus on an intuitive cost-driven state representation learning method for solving Linear Quadratic Gaussian (LQG) control, one of the most fundamental partially observable control problems. As our main results, we establish finite-sample guarantees of finding a near-optimal state representation function and a near-optimal controller using the directly learned latent model. To the best of our knowledge, despite various empirical successes, prior to this work it was unclear if such a cost-driven latent model learner enjoys finite-sample guarantees. Our work underscores the value of predicting multi-step costs, an idea that is key to our theory, and notably also an idea that is known to be empirically valuable for learning state representations.
@@ 1 Introduction
Various empirical attempts (Schrittwieser et al., 2020; Zhang et al., 2020; Okada and Taniguchi, 2021; Deng et al., 2021; Yang et al., 2022) have been made to bypass observation reconstruction. Apart from observation, the interaction involves two other variables: actions (control inputs) and costs. Inverse model methods (Lamb et al., 2022) reconstruct actions; while other methods rely on costs. We argue that since neither the reconstruction function nor the inverse model is used for policy learning, cost-driven state representation learning is the most direct one, in that costs are directly relevant for control purposes. In this paper, we aim to examine the soundness of this methodology in linear quadratic Gaussian (LQG) control, one of the most fundamental partially observable control models.
Parallel to the empirical advances of learning for control from pixels, partially observable linear systems has been extensively studied in the context of learning for dynamic control (Oymak and Ozay, 2019; Simchowitz et al., 2020; Lale et al., 2020, 2021; Zheng et al., 2021; Minasyan et al., 2021; Umenberger et al., 2022). In this context, the representation function is more formally referred to as a filter, the optimal one being the Kalman filter. Most existing model-based learning approaches for LQG control focus on the linear time-invariant (LTI) case, and are based on the idea of learning Markov parameters (Ljung, 1998), the mapping from control inputs to observations. Hence, they need to predict observations by definition. Motivated by the empirical successes in control from pixels, we take a different, cost-driven route, in hope of avoiding reconstructing observations or control inputs.
@@
Can direct, cost-driven state representation learning provably solve LQG control?
@@
This work answers the question in the affirmative, by establishing finite-sample guarantees for for a cost-driven state representation learning method.
@@
Challenges & Our techniques. Overall, to establish finite-sample guarantees, a major technical challenge is to deal with the quadratic regression problem in cost prediction, arising from the inherent quadratic form of the cost function in LQG. Directly solving the problem for the representation function involves quartic optimization; instead, we propose to solve a quadratic regression problem, followed by low-rank approximate factorization. The quadratic regression problem also appears in identifying the cost matrices, which involves concentration for random variables that are fourth powers of Gaussians. We believe these techniques might be of independent interest.
@@
Moreover, the first $\ell$-step latent states may not be adequately excited (having full-rank covariance), which invalidates the use of most system identification techniques. We instead identify only relevant directions of the system parameters, and prove that this is sufficient for learning a near-optimal controller by analyzing state covariance mismatch. This fact is reflected in the separation in the statement of Theorem 1; developing finite-sample analysis in this case is technically challenging.
@@
Implications. For practitioners, one takeaway from our work is the benefit of predicting multi-step cumulative costs in cost-driven state representation learning. Whereas the cost at a single time step may not be revealing enough of the latent state, the cumulative cost across multiple steps can be. This is an intuitive idea for the control community, given the multi-step nature in the classical definitions of controllability and observability. Its effectiveness has also been empirically observed in MuZero (Schrittwieser et al., 2020) in state representation learning for control, and our work can be viewed as a formal understanding of it in the LQG control setting.
@@
Notation. We use 00 (resp. $1$) to denote either the scalar or a matrix consisting of all zeros (resp. all ones); we use $I$ to denote an identity matrix. The dimension, when emphasized, is specified in subscripts, e.g., $0_{d_{x}\times d_{x}},1_{d_{x}},I_{d_{x}}$. Let $\mathbb{I}_{S}$ denote the indicator function for set $S$ and $\mathbb{I}_{S}(A)$ apply to matrix $A$ elementwise. For some positive semidefinite $P$, we define $\|v\|_{P}:=(v^{\top}Pv)^{1/2}$. Semicolon “;” denotes stacking vectors or matrices vertically. For a collection of $d$-dimensional vectors $(v_{t})_{t=i}^{j}$, let $v_{i:j}:=[v_{i};v_{i+1};\ldots;v_{j}]\in\mathbb{R}^{d(j-i+1)}$ denote the concatenation along the column. For random variable $\eta$, let $\|\eta\|_{\psi_{\beta}}$ denote its $\beta$-sub-Weibull norm, a special case of Orlicz norms (Zhang and Wei, 2022), with $\beta=1,2$ corresponding to subexponential and sub-Gaussian norms. $\sigma_{i}(A),\sigma_{\min}(A),\sigma_{\min}^{+}(A),\sigma_{\max}(A)$ denote its $i$th largest, minimum, minimum positive, maximum singular values, respectively. $\|A\|_{2},\|A\|_{F},\|A\|_{\ast}$ denote the operator (induced by vector $2$-norms), Frobenius, nuclear norms of matrix $A$, respectively. $\left\langle\cdot,\cdot\right\rangle_{F}$ denotes the Frobenius inner product between matrices. The Kronecker, symmetric Kronecker and Hadamard products between matrices are denoted by $\otimes$”, “$\otimes_{s}$” and $\odot$”, respectively. $\mathrm{vec}(\cdot)$ and $\mathrm{svec}(\cdot)$ denote flattening a matrix and a symmetric matrix by stacking their columns; $\mathrm{svec}(\cdot)$ does not repeat the off-diagonal elements, but scales them by $\sqrt{2}$ (Schacke, 2004). Let $\textnormal{diag}(\cdot)$ denote the block diagonal matrix formed by the matrices inside the parentheses.
@@ 2 Problem setup
We study a partially observable linear time-varying (LTV) dynamical system
@@
Assumption 1 is standard in controlling LTV systems (Zhou and Zhao, 2017; Minasyan et al., 2021), satisfied by a stable LTI system. It essentially says that zero control is a stabilizing policy, and can be potentially relaxed to the assumption of being given a stabilizing policy (Lale et al., 2020; Simchowitz et al., 2020). Specifically, one can excite the system using the stabilizing policy plus Gaussian random noises.
@@
for $\ell-1\leq t\leq T-1$, which reduces to the standard controllability matrix $[B,\ldots,A^{\ell-1}B]$ in the LTI setting. We make the following controllability assumption.
@@
so Assumption 2 ensures that from any state $x$, there exist control inputs that drive the state to 00 in $\ell$ steps, and $\nu$ ensures that the equation leading to them is well conditioned. We do not assume controllability for $10\leq t<\ell-$ 1, since we do not want to impose the constraint that $d_{u}>d_{x}$. This turns out to present a significant challenge for state representation learning, as seen from the separation of the results before and after the $\ell$-steps in Theorem 1.
@@
$c_{t}(x,u)=\|x\|_{Q^{\ast}_{t}}^{2}+\|u\|_{R^{\ast}_{t}}^{2},\quad 0\leq t\leq T-1,$ (2.2)
and $c_{T}(x)=\|x\|_{Q^{\ast}_{T}}^{2}$, where $(Q^{\ast}_{t})_{t=0}^{T}$ are positive semidefinite matrices and $(R^{\ast}_{t})_{t=0}^{T-1}$ are positive definite matrices. Sometimes the cost is defined as a function on observation $y$. Since the quadratic form $y^{\top}Q^{\ast}_{t}y=x^{\top}(C^{\ast}_{t})^{\top}Q^{\ast}_{t}C^{\ast}_{t}x$, our analysis still applies if the assumptions on $(Q^{\ast}_{t})_{t=0}^{T}$ hold for $((C^{\ast}_{t})^{\top}Q^{\ast}_{t}C^{\ast}_{t})_{t=0}^{T}$ instead.
$(A,C)$ and $(A,Q^{1/2})$ observabilities are standard assumptions in controlling LTI systems. To differentiate from the former, we call the latter cost observability, since it implies the states are observable through costs. Whereas Markov-parameter-based approaches need to assume $(A,C)$ observability to identify the system, our cost-driven approach does not. Here we deal with the more difficult problem of having only the scalar cost as the supervision signal (instead of the concatenation of all observations, as in Markov-parameter-based ones). Nevertheless, the notion of cost observability is still important for our approach, formally defined as follows.
@@
For all $10\leq t\leq\ell-$ 1, $Q^{\ast}_{t}\succcurlyeq\mu^{2}I$. For all $\ell\leq t\leq T$, there exists $m>0$ such that the cost observability Gramian (Kailath, 1980)
@@
This assumption ensures that without noises, if we start with a nonzero state, the cumulative cost becomes positive in $m$ steps. The special requirement for $10\leq t\leq\ell-$ 1 results from the difficulty in lacking controllability in these time steps. The following is a regularity assumption on system parameters.
@@
$(\sigma_{\min}(\Sigma_{v_{t}}))_{t=0}^{T}$ are uniformly lower bounded by some $\sigma_{v}>0$. The operator norms of all matrices in the problem definition are uniformly upper bounded, including $(A^{\ast}_{t},B^{\ast}_{t},R^{\ast}_{t},\Sigma_{w_{t}})_{t=0}^{T-1}$, $(C^{\ast}_{t},Q^{\ast}_{t},\Sigma_{v_{t}})_{t=0}^{T}$. In other words, they are all $\mathcal{O}(1)$.
Let $h_{t}:=[y_{0:t};u_{0:(t-1)}]\in\mathbb{R}^{(t+1)d_{y}+td_{u}}$ denote the available history before deciding control $u_{t}$. A policy $\pi=(\pi_{t}:h_{t}\mapsto u_{t})_{t=0}^{T-1}$ determines at time $t$ a control input $u_{t}$ based on history $h_{t}$. With a slight abuse of notation, let $c_{t}:=c_{t}(x_{t},u_{t})$ for $10\leq t\leq T-$ 1 and $c_{T}:=c_{T}(x_{T})$ denote the cost at each time step. Then, $J^{\pi}:=\mathbb{E}^{\pi}[\sum_{t=0}^{T}c_{t}]$ is the expected cumulative cost under policy $\pi$, where the expectation is taken over the randomness in the process noises, observation noises, and controls (if the policy $\pi$ is stochastic). The objective of LQG control is to find a policy $\pi$ such that $J^{\pi}$ is minimized.
If the system parameters $((A^{\ast}_{t},B^{\ast}_{t},R^{\ast}_{t})_{t=0}^{T-1},(C^{\ast}_{t},Q^{\ast}_{t})_{t=0}^{T})$ are known, the optimal control is obtained by combining the Kalman filter $z^{\ast}_{0}=L^{\ast}_{0}y_{0}$,
@@
for $10\leq t\leq T-$ 1, with the optimal feedback control gains of the linear quadratic regulator (LQR) $(K^{\ast}_{t})_{t=0}^{T-1}$ such that $u^{\ast}_{t}=K^{\ast}_{t}z^{\ast}_{t}$, where $(L^{\ast}_{t})_{t=0}^{T}$ are the Kalman gains; this is known as the separation principle (Åström, 2012). The Kalman gains and optimal feedback control gains are given by
$L^{\ast}_{t}=S^{\ast}_{t}(C^{\ast}_{t})^{\top}(C^{\ast}_{t}S^{\ast}_{t}(C^{\ast}_{t})^{\top}+\Sigma_{v_{t}})^{-1},$
$K^{\ast}_{t}=-((B^{\ast}_{t})^{\top}P^{\ast}_{t+1}B^{\ast}_{t}+R_{t})^{-1}(B^{\ast}_{t})^{\top}P^{\ast}_{t+1}A^{\ast}_{t},$
@@
$S^{\ast}_{t+1}=A^{\ast}_{t}(S^{\ast}_{t}-S^{\ast}_{t}(C^{\ast}_{t})^{\top}(C^{\ast}_{t}S^{\ast}_{t}(C^{\ast}_{t})^{\top}+\Sigma_{v_{t}})^{-1}C^{\ast}_{t}S^{\ast}_{t})(A^{\ast}_{t})^{\top}+\Sigma_{w_{t}},$ (2.3)
$P^{\ast}_{t}=(A^{\ast}_{t})^{\top}(P^{\ast}_{t+1}-P^{\ast}_{t+1}B^{\ast}_{t}((B^{\ast}_{t})^{\top}P^{\ast}_{t+1}B^{\ast}_{t}+R^{\ast}_{t})^{-1}(B^{\ast}_{t})^{\top}P^{\ast}_{t+1})A^{\ast}_{t}+Q^{\ast}_{t},$ (2.4)
@@
We consider data-driven control in a partially observable LTV system (2.1) with unknown cost matrices $(Q^{\ast}_{t})_{t=0}^{T}$. For simplicity, we assume $(R^{\ast}_{t})_{t=0}^{T}$ is known, though our approaches can be readily generalized to the case without knowing them; one can identify them in the quadratic regression (3.3).
2.1 Latent model of LQG
Under the Kalman filter, the observation prediction error $i_{t+1}:=y_{t+1}-C^{\ast}_{t+1}(A^{\ast}_{t}z^{\ast}_{t}+B^{\ast}_{t}u_{t})$ is called an innovation. It is known that $i_{t}$ is independent of history $h_{t}$ and $(i_{t})_{t=1}^{T}$ are independent (Bertsekas, 2012).
Now we are ready to present the following proposition that represents the system in terms of the state estimates by the Kalman filter, which we shall refer to as the latent model.

### v3

# Cost-Driven Representation Learning for Linear Quadratic Gaussian Control: Part I

@@ Abstract
We study the task of learning state representations from potentially high-dimensional observations, with the goal of controlling an unknown partially observable system. We pursue a cost-driven approach, where a dynamic model in some latent state space is learned by predicting the costs without predicting the observations or actions. In particular, we focus on an intuitive cost-driven state representation learning method for solving Linear Quadratic Gaussian (LQG) control, one of the most fundamental partially observable control problems. As our main results, we establish finite-sample guarantees of finding a near-optimal state representation function and a near-optimal controller using the directly learned latent model, for finite-horizon time-varying LQG control problems. To the best of our knowledge, despite various empirical successes, finite-sample guarantees of such a cost-driven approach remain elusive. Our result underscores the value of predicting multi-step costs, an idea that is key to our theory, and notably also an idea that is known to be empirically valuable for learning state representations. A second part of this work, that is to appear as Part II, addresses the infinite-horizon linear time-invariant setting; it also extends the results to an approach that implicitly learns the latent dynamics, inspired by the recent empirical breakthrough of MuZero in model-based reinforcement learning.
@@ 1 Introduction
Various empirical attempts (Schrittwieser et al., 2020; Zhang et al., 2020; Okada and Taniguchi, 2021; Deng et al., 2021; Yang et al., 2022) have been made to bypass observation reconstruction. Apart from observation, the interaction involves two other variables: actions (control inputs) and costs. Inverse model methods (Lamb et al., 2022) reconstruct actions; while other methods rely on costs. We argue that since neither the reconstruction function nor the inverse model is used for policy learning, cost-driven state representation learning is the most direct one, in that costs are directly relevant for control purposes. In this work, we aim to examine the soundness of this methodology in linear quadratic Gaussian (LQG) control, one of the most fundamental partially observable control models.
Parallel to the empirical advances of learning for control from pixels, partially observable linear systems has been extensively studied in the context of learning for dynamic control (Oymak and Ozay, 2019; Simchowitz et al., 2020; Lale et al., 2020, 2021; Zheng et al., 2021; Minasyan et al., 2021; Umenberger et al., 2022). In this context, the state representation function is more formally referred to as a filter, the optimal one being the Kalman filter. Most existing model-based learning approaches for LQG control focus on the linear time-invariant (LTI) case, and are based on the idea of learning Markov parameters (Ljung, 1998), the mapping from control inputs to observations. Hence, they need to predict observations by definition. Motivated by the empirical successes in control from pixels, we take a different, cost-driven route, in hope of avoiding reconstructing observations or control inputs.
@@
Can cost-driven state representation learning provably solve LQG control?
@@
This work answers the question in the affirmative, by establishing finite-sample guarantees for a cost-driven state representation learning method. We address the finite-horizon linear time-varying (LTV) setting in this Part I, and will move on to the infinite-horizon linear time-invariant (LTI) setting with additional technical challenges in Part II of the work.
@@
Challenges & Our techniques. To establish finite-sample guarantees, a major technical challenge is to deal with the quadratic regression problem in cost prediction, arising from the inherent quadratic form of the LQG cost. Directly solving for the state representation function involves quartic optimization; instead, we propose to solve a quadratic regression problem, followed by low-rank approximate factorization. The quadratic regression problem also appears in identifying the cost matrices, involving the concentration for random variables that are fourth powers of Gaussians. Our techniques to address these challenges may be of independent interest.
@@
Moreover, the first $\ell$-step latent states may not be adequately excited (with full-rank covariance), which results in the identification of the latent model only in partial directions. This poses a significant challenge for certifying the performance of the learned controller, as the learned latent model from which we synthesize the controller may be neither stable nor controllable. We overcome this challenge by analyzing state covariance mismatch using induction, showing that identifying only the relevant directions suffices for learning a near-optimal controller. This fact is reflected in the dependence on $\ell$ in the statement of Theorem 1.
Lastly, the learned latent states and the errors in the learned latent states are correlated as they are both functions of the same observed trajectory. This challenge arises both in analyzing latent model identification and in certifying the performance of the learned controller. We tackle this challenge by modeling the errors as general correlated perturbations whose magnitudes are controlled by the errors in the learned state representation function.
@@
Implications. For practitioners, one takeaway is the benefit of predicting multi-step cumulative costs in cost-driven state representation learning. Whereas the cost at a single time step may not be revealing enough of the latent state, the cumulative cost across multiple steps can be. This is an intuitive idea for the control community, given the multi-step nature in the classical definitions of controllability and observability. Its effectiveness has also been empirically observed in MuZero (Schrittwieser et al., 2020) in state representation learning for control, and our work can be viewed as a formal understanding of it in the LQG setting.
@@
Notation. We use 0 (resp. $1$) to denote either the scalar or a matrix consisting of all zeros (resp. all ones); we use $I$ to denote an identity matrix. The dimension, when emphasized, is specified in subscripts, e.g., $0_{d_{x}\times d_{x}},1_{d_{x}},I_{d_{x}}$. Let $\mathbb{I}_{S}$ denote the indicator function for set SS and $\mathbb{I}_{S}(A)$ apply to matrix $A$ elementwise. For some positive semidefinite $P$, we define $\|v\|_{P}:=(v^{\top}Pv)^{1/2}$. Semicolon “;” denotes stacking vectors or matrices vertically. For a collection of $d$-dimensional vectors $(v_{t})_{t=i}^{j}$, let $v_{i:j}:=[v_{i};v_{i+1};\ldots;v_{j}]\in\mathbb{R}^{d(j-i+1)}$ denote the concatenation along the column. For random variable $\eta$, let $\|\eta\|_{\psi_{\beta}}$ denote its $\beta$-sub-Weibull norm, a special case of Orlicz norms (Zhang and Wei, 2022), with $\beta=1,2$ corresponding to subexponential and sub-Gaussian norms. For matrix $A$, let $\sigma_{i}(A),\sigma_{\min}(A),\sigma_{\min}^{+}(A),\sigma_{\max}(A)$ denote its $i$th largest, minimum, minimum positive, maximum singular values, respectively. $\|A\|_{2},\|A\|_{F},\|A\|_{\ast}$ denote the operator (induced by vector $2$-norms), Frobenius, nuclear norms of matrix $A$, respectively. $\left\langle\cdot,\cdot\right\rangle_{F}$ denotes the Frobenius inner product between matrices. For square matrix $A$, let $\lambda_{\min}(A)$ be its minimum eigenvalue. The Kronecker, symmetric Kronecker, and Hadamard products between matrices are denoted by $\otimes$”, $\otimes_{s}$” and $\odot$”, respectively. $\mathrm{vec}(\cdot)$ and $\mathrm{svec}(\cdot)$ denote flattening a matrix and a symmetric matrix by stacking their columns; $\mathrm{svec}(\cdot)$ does not repeat the off-diagonal elements, but scales them by $\sqrt{2}$ (Schacke, 2004). Let $\textnormal{diag}(\cdot)$ denote the block diagonal matrix formed by the matrices inside the parentheses. For $\Sigma_{i=a}^{b}$, we define the sum to be zero if the lower index $a$ is greater than the upper index $b$.
@@ 2 Problem setup
We study a partially observable linear dynamical system
@@
Assumption 1 is standard in controlling LTV systems (Zhou and Zhao, 2017; Hu and Mitsui, 2017; Minasyan et al., 2021), satisfied by a stable LTI system. It essentially says that zero control is a stabilizing controller, and can be potentially relaxed to the assumption of being given a stabilizing controller as in Lale et al. (2020); Simchowitz et al. (2020), where one can excite the system using the stabilizing controller plus Gaussian random noises.
@@
for $\ell-1\leq t\leq T-1$, which reduces to the standard controllability matrix $B,\ldots,A^{\ell-1}B]$ in the LTI setting. We make the following controllability assumption.
@@
so Assumption 2 ensures that from any state $x$, there exist control inputs that drive the state to 0 in $\ell$ steps, and $\nu$ ensures that the equation leading to them is well conditioned. We do not assume controllability for $0\leq t<\ell-1$, since we do not want to impose the constraint that $d_{u}>d_{x}$. This turns out to present a significant challenge for analyzing state representation function learning, latent model identification, and the performance of the overall policy, resulting in the separation at step $\ell$ in the sample complexity guarantees (see Theorem 1).
@@
$c_{t}(x,u)=\|x\|_{Q^{\ast}_{t}}^{2}+\|u\|_{R^{\ast}_{t}}^{2},\quad 0\leq t\leq T-1,\quad c_{T}(x)=\|x\|_{Q^{\ast}_{T}}^{2},$ (2.2)
where $(Q^{\ast}_{t})_{t=0}^{T}$ are positive semidefinite matrices and $(R^{\ast}_{t})_{t=0}^{T-1}$ are positive definite matrices. Sometimes the cost is defined as a function of the observation $y$. Since the quadratic form $y^{\top}Q^{\ast}_{t}y=x^{\top}(C^{\ast}_{t})^{\top}Q^{\ast}_{t}C^{\ast}_{t}x$, our analysis still applies if the assumptions on $(Q^{\ast}_{t})_{t=0}^{T}$ hold for $((C^{\ast}_{t})^{\top}Q^{\ast}_{t}C^{\ast}_{t})_{t=0}^{T}$ instead.
The observability assumptions on $(A,C)$ and $(A,Q^{1/2})$ are standard in controlling LTI systems. To differentiate from the former, we call the latter cost observability, since it implies the states are observable through costs. Whereas Markov-parameter-based approaches need to assume $(A,C)$ observability to identify the system, our cost-driven approach does not. Here we deal with the more difficult problem of having only the scalar cost as the supervision signal (instead of the concatenation of all observations, as in Markov-parameter-based ones). Nevertheless, the notion of cost observability is still important for our approach, formally defined as follows.
@@
For all $0\leq t\leq\ell-1$, $Q^{\ast}_{t}\succcurlyeq\mu^{2}I$. For all $\ell\leq t\leq T$, there exists $m>0$ such that the cost observability Gram matrix (Kailath, 1980)
@@
This assumption ensures that without noises, if we start with a nonzero state, the cumulative cost becomes positive in $m$ steps. The special requirement for $0\leq t\leq\ell-1$ results from the difficulty in lacking controllability in these time steps. The following is a regularity assumption on system parameters.
@@
$(\lambda_{\min}(\Sigma_{v_{t}}))_{t=0}^{T}$ and $(\lambda_{\min}(R^{\ast}_{t}))_{t=0}^{T-1}$ are uniformly lower bounded; that is, they are all $\Omega(1)$. The operator norms of all matrices in the problem definition and $(M^{\ast}_{t})_{t=0}^{T}$ to be defined in §2.1 are uniformly upper bounded, including $(A^{\ast}_{t},B^{\ast}_{t},R^{\ast}_{t},\Sigma_{w_{t}})_{t=0}^{T-1}$, $(C^{\ast}_{t},Q^{\ast}_{t},\Sigma_{v_{t}})_{t=0}^{T}$; that is, they are all $\mathcal{O}(1)$.
Along with Assumption 1, this assumption ensures that under control $u_{t}\sim\mathcal{N}(0,\sigma_{u}^{2}I)$ for all $t\geq 0$ with $\sigma_{u}=\mathcal{O}(1)$, the covariance matrices $\mathrm{Cov}(x_{t})$ and $\mathrm{Cov}(y_{t})$ have operator norms bounded by $\mathcal{O}(1)$ for all $t\geq 0$. Moreover, $(S^{\ast}_{t},L^{\ast}_{t},P^{\ast}_{t})_{t=0}^{T}$ and $(K^{\ast}_{t})_{t=0}^{T-1}$, to be defined shortly in (2.3) to (2.6) in the optimal solution, have $\mathcal{O}(1)$ operator norms.
Let $h_{t}:=[y_{0:t};u_{0:(t-1)}]\in\mathbb{R}^{(t+1)d_{y}+td_{u}}$ denote the available history before deciding control $u_{t}$ for $t\geq 1$ and define $h_{0}:=y_{0}$. A policy $\pi=(\pi_{t}:h_{t}\mapsto u_{t})_{t=0}^{T-1}$ determines at time $t$ a control input $u_{t}$ based on history $h_{t}$. With a slight abuse of notation, let $c_{t}:=c_{t}(x_{t},u_{t})$ for $0\leq t\leq T-1$ and $c_{T}:=c_{T}(x_{T})$ denote the cost at each time step. Then, $J(\pi):=\mathbb{E}^{\pi}[\sum_{t=0}^{T}c_{t}]$ is the expected cumulative cost under policy $\pi$, where the expectation is taken over the randomness in the process noises, observation noises, and controls (if the policy $\pi$ is stochastic). The objective of LQG control is to find a policy $\pi$ such that $J(\pi)$ is minimized.
If the system parameters $((A^{\ast}_{t},B^{\ast}_{t},R^{\ast}_{t})_{t=0}^{T-1},(C^{\ast}_{t},Q^{\ast}_{t})_{t=0}^{T})$ are known, optimal control is obtained by combining the Kalman filter $z^{\ast}_{0}=L^{\ast}_{0}y_{0}$,
@@
for $0\leq t\leq T-1$, with the optimal feedback control gains of the linear quadratic regulator (LQR) $(K^{\ast}_{t})_{t=0}^{T-1}$ such that $u^{\ast}_{t}=K^{\ast}_{t}z^{\ast}_{t}$, where $(L^{\ast}_{t})_{t=0}^{T}$ are the Kalman gains; this is known as the separation principle (Åström, 2012). The Kalman gains and optimal feedback control gains are given by
$L^{\ast}_{t}=S^{\ast}_{t}(C^{\ast}_{t})^{\top}(C^{\ast}_{t}S^{\ast}_{t}(C^{\ast}_{t})^{\top}+\Sigma_{v_{t}})^{-1},$ (2.3)
$K^{\ast}_{t}=-((B^{\ast}_{t})^{\top}P^{\ast}_{t+1}B^{\ast}_{t}+R^{\ast}_{t})^{-1}(B^{\ast}_{t})^{\top}P^{\ast}_{t+1}A^{\ast}_{t},$ (2.4)
@@
$\begin{aligned} S^{\ast}_{t+1}=\;&A^{\ast}_{t}(S^{\ast}_{t}-S^{\ast}_{t}(C^{\ast}_{t})^{\top}(C^{\ast}_{t}S^{\ast}_{t}(C^{\ast}_{t})^{\top}+\Sigma_{v_{t}})^{-1}C^{\ast}_{t}S^{\ast}_{t})(A^{\ast}_{t})^{\top}+\Sigma_{w_{t}},\end{aligned}$ (2.5)
$\begin{aligned} P^{\ast}_{t}=\;&(A^{\ast}_{t})^{\top}(P^{\ast}_{t+1}-P^{\ast}_{t+1}B^{\ast}_{t}((B^{\ast}_{t})^{\top}P^{\ast}_{t+1}B^{\ast}_{t}+R^{\ast}_{t})^{-1}(B^{\ast}_{t})^{\top}P^{\ast}_{t+1})A^{\ast}_{t}+Q^{\ast}_{t},\end{aligned}$ (2.6)
@@
We consider data-driven control in a partially observable LTV system (2.1) with unknown cost matrices $(Q^{\ast}_{t})_{t=0}^{T}$. For simplicity, we assume $(R^{\ast}_{t})_{t=0}^{T}$ is known, though our approaches can be readily generalized to the case where they are unknown; one can identify them in the quadratic regression (3.3).
2.1 Latent model of finite-horizon time-varying LQG
Under the Kalman filter, the observation prediction error $i_{t+1}:=y_{t+1}-C^{\ast}_{t+1}(A^{\ast}_{t}z^{\ast}_{t}+B^{\ast}_{t}u_{t})$ is called an innovation. It is known that $i_{t}$ is independent of history $h_{t}$ and $(i_{t})_{t=1}^{T}$ are independent (Bertsekas, 2012). Now we are ready to present the following proposition that represents the system in terms of the state estimates by the Kalman filter, which we shall refer to as the latent model.